• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 16:03:49 on June26,2000

    Entry number 44919

    keyword=Shift Summary, 0800-1600, June 26, 2000




    Shift Summary, 0800-1600, June 26, 2000
    ___________________________________

    0800 Still no beam. MCC (according to ROM) predicts beam may return by noon.
    All experimental equipment functioning fine, except FPP has noise.

    1000-1240 Jones+Kelly enter hall to investigate FPP noise problems. Details
    in separate halog by Jones.

    1515: see some 80 uA beam, but jumping around
    1525: magnet problems near injector.
    1535: start a run! Run #2547, 80 uA.
    1600: FPP still has problems. Requested access.